# What Breaks Knowledge Graph based RAG? Empirical Insights into Reasoning under Incomplete Knowledge

### v3

@@ Abstract
Knowledge Graph-based Retrieval-Augmented Generation (KG-RAG) is an increasingly explored approach for combining the reasoning capabilities of large language models with the structured evidence of knowledge graphs. However, current evaluation practices fall short: existing benchmarks often include questions that can be directly answered using existing triples in KG, making it unclear whether models perform reasoning or simply retrieve answers directly. Moreover, inconsistent evaluation metrics and lenient answer matching criteria further obscure meaningful comparisons. In this work, we introduce a general method for constructing benchmarks, together with an evaluation protocol, to systematically assess KG-RAG methods under knowledge incompleteness. Our empirical results show that current KG-RAG methods have limited reasoning ability under missing knowledge, often rely on internal memorization, and exhibit varying degrees of generalization depending on their design.
What Breaks Knowledge Graph based RAG? Empirical Insights
into Reasoning under Incomplete Knowledge
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has become a widely adopted framework for enhancing large language models (LLMs) by incorporating external knowledge through a retrieve-then-generate paradigm (Khandelwal2020rag; Izacard2021FiD; Borgeaud2022retro; ram2023ICRALM; zhu2025argrag). By conditioning the generation on the retrieved documents, RAG enables LLMs to answer questions or perform tasks using more comprehensive and up-to-date knowledge than what is stored in their parameters. To improve retrieval accuracy, enable structured reasoning and support explanation, recent research has pivoted toward RAG methods based on Knowledge Graph (KG-RAG) (han2024retrieval; peng2024graph). Most of them directly use existing knowledge graphs (KGs) (linhao2024rog; Jiashuo2024ToG; zhou2025gr), while some of them construct and extend structured knowledge from unstructured documents (fang2024reano). Such KG-RAG systems are expected to be well-suited for structured reasoning, where answers require synthesizing information from multiple connected facts.
Despite growing interest, current evaluation practices for KG-RAG fall short in two key ways. First, most existing benchmarks  (Yih2016webqsp; Talmor2018WSQ) are constructed on top of complete KGs, where direct evidence supporting the answer is readily available. For example, given the question “Who is the brother of Justin Bieber?”, the KG contains the triple hasBrother(JustinBieber, JaxonBieber), allowing the system to answer the question without reasoning. However, real-world KGs are often incomplete, and answering such questions in practice may require reasoning over alternative paths, e.g., combining hasParent(JustinBieber, JeremyBieber) and hasChild(JeremyBieber, JaxonBieber) to infer the sibling relationship. As a result, current benchmarks do not assess whether KG-RAG methods can reason over missing knowledge or simply retrieve answers directly from explicit evidence.
@@
In this work, we present a general method for constructing benchmarks and evaluation protocols to systematically assess KG-RAG methods under knowledge incompleteness. Each question in our benchmark is constructed such that it cannot be answered using a single explicit triple. Instead, the answer must be inferred by reasoning over alternative paths in the KG. To construct this benchmark, we follow a two-step process: (1) we mine high-confidence logical rules from the KG using a rule mining algorithm, and (2) we generate natural language questions based on rule groundings, ensuring that the directly supporting triple is removed while the remaining KG still contains sufficient evidence to infer the answer.
Our empirical study on this benchmark reveals several key limitations of current KG-RAG systems. First, most models struggle to recover answers when direct supporting facts are removed, highlighting their limited reasoning capacity. Second, training-based methods (e.g., RoG, GNN-RAG) show stronger robustness under KG incompleteness compared to non-trained systems. Third, we find that textual entity labels substantially boost performance, suggesting that LLMs rely heavily on internal memorization.
@@ 2.1 Knowledge Graph Question Answering (KGQA)
Approaches to KGQA can be broadly categorized into the following two types: (1) Semantic parsing-based methods Yih2016webqsp; gu2021beyond; ye2021rng translate questions into formal executable queries (e.g., SPARQL or logical forms), which are then executed over the KG to retrieve the answer entities. These methods offer high precision and interpretability, as the reasoning process is explicitly encoded. However, they face several challenges, including difficulty in understanding semantically and syntactically complex questions, handling diverse and compositional logical forms, and managing the large search space involved in parsing multi-relation querieslan2021survey. (2) Embedding-based methods yao2019kg; baek2023direct, by contrast, encode questions and entities into a shared vector space and rank answer candidates based on embedding similarity. While these methods are end-to-end trainable and do not require annotated logical forms, they often struggle with multi-hop reasoning qiu2020stepwise, lack interpretability biswas2023knowledge, and exhibit high uncertainty in their predictions zhu-etal-2024-predictive; zhu2025conformalized; zhu-etal-2025-predicate; zhu2025certainty.
@@ 2.2 KGQA with RAG
Recently, a new line of work has emerged that integrates large language models more tightly into the KG reasoning process. These KG-RAG methods go beyond previous KGQA approaches by coupling structured retrieval with generative reasoning capabilities of LLMs. RoG linhao2024rog adopts a planning-retrieval-reasoning pipeline, where an LLM generates relation paths as plans, retrieves corresponding paths from the KG, and reasons over them for answer generation. G-Retriever he2024g retrieves subgraphs via Prize-Collecting Steiner Tree optimization and provides them to LLMs as soft prompts or text prompts for question answering. GNN-RAG mavromatis2024gnn uses a GNN to select candidate answers and retrieves shortest paths to them, which are verbalized for LLM-based answer generation. PoG chen2024pog instead decomposes questions into subgoals and iteratively explores and self-corrects reasoning paths using a guidance-memory-reflection mechanism. ToG Jiashuo2024ToG treats the LLM as an agent that interactively explores and aggregates evidence on KGs through iterative beam search, StructGPT jiang2023structgpt introduces an interface-based framework that enables LLMs to iteratively extract relevant evidence from structured data such as KGs and reason over the retrieved information to answer complex questions.
@@ 2.3 KGQA Benchmarks
Existing benchmarks such as WebQSP Yih2016webqsp, CWQ talmor2018web, and GrailQA gu2021beyond are widely used to evaluate KGQA. However, during dataset construction, only question-answer pairs that yield a gold SPARQL answer on the reference KG are retained, while any unanswerable questions are discarded Yih2016webqsp; talmor2018web; gu2021beyond. This implicitly assumes that each question can be answered directly using an existing fact in the KG, overlooking the reality that KGs are often incomplete. To address this, recent works xu2024generate; zhou2025evaluating simulate KG incompleteness by either randomly deleting triples from the whole KG or removing those along the shortest path(s) between the question and answer entities. However, these approaches have a key limitation: they cannot ensure that sufficient knowledge remains in the KG to support answering each question. This can lead to misleading evaluations, as performance drops may stem from unanswerable questions rather than limitations in the model’s reasoning ability.
@@ 2.4 LLM Reasoning Evaluation
Evaluation of LLM reasoning typically covers both logical question answering and logical consistency, with benchmarks spanning deduction, multiple-choice, proof generation, and consistency constraints such as entailment, negation, and transitivity. Datasets like LogicBench parmar2024logicbench, ProofWriter tafjord2020proofwriter, and LogicNLI tian2021diagnosing are widely used for these tasks, but results indicate that LLMs still struggle with robust logical reasoning cheng2025empowering; he2025supposedly.
@@ 3.1 Rule Mining
To ensure that questions in our benchmark require reasoning rather than direct lookup, we first identify triples that are logically inferable from other facts. We achieve this by mining high-confidence Horn rules from the original KG using the AMIE3 algorithm (lajus2020fast).
AMIE3 is a widely used rule mining system designed to operate efficiently over large-scale KGs. A logical rule discovered by AMIE3 has the following form (Horn rules horn1951sentences):
@@
We retain only rules with high confidence and sufficient support, filtering out noisy or spurious patterns. Specifically, we run AMIE3 with a confidence threshold of 0.3, a head coverage threshold of 0.1, and a maximum rule length of 4. AMIE3 incrementally generates candidate rules via breadth-first refinement lajus2020fast and evaluates them using confidence and head coverage; only those meeting the specified thresholds are retained. Additional details on the rule generation and filtering process are provided in Appendix A.
@@ 3.2 Dataset Generation
KGs typically exhibit a "long-tail" distribution, where a small number of entities participate in a disproportionately large number of triples, while the majority appear only infrequently  (mohamed2020popularity; chen2023knowledge). This imbalance can cause many generated questions to share the same answer entity, leading to biased evaluation.
@@ 3.3 Dataset Overview
To support a systematic evaluation of reasoning under knowledge incompleteness, we construct benchmark datasets based on three well established KGs: Family sadeghian2019drum, FB15k-237 toutanova2015observed and Wikidata5m wang2021kepler. These datasets differ in size, structure, and domain coverage, enabling evaluation across both synthetic and real-world settings.
@@
  • FB15k-237 is a widely used benchmark derived from Freebase dong2014data. The graph spans 14,541 entities and 237 predicates, covering real-world domains such as people, locations, and organizations.
  • Wikidata5m is a large-scale real-world KG constructed from Wikidata (vrandevcic2014wikidata). It contains about 4.6 million entities, 822 relations, and over 20 million triples spanning diverse domains such as people, places, organizations, and events wang2021kepler.
@@
Table 2: Examples from our benchmark datasets. Each instance includes a natural-language question, a topic entity (provided to the KG-RAG model), and the full set of correct answers. The red-highlighted answer denotes the hard answer—i.e., the one whose supporting triple has been removed in the incomplete KG setting. We also show the corresponding direct evidence (the deleted triple) and an alternative path derived from a mined rule that enables inference of the answer.
Datasets.
@@
Table 3: Dataset statistics.
@@ 4.1 Standardization Gaps in Prior Work
Many studies rely on vague textual descriptions such as “whether the top-1 predicted answer is correct” for Hits@1 without specifying critical details (linhao2024rog; mavromatis2024gnn; he2024g). It is often unclear how LLM outputs are parsed, how “top-1” is defined in the absence of explicit ranking, and what constitutes correctness (e.g., exact match vs. substring inclusion). Some works omit definitions entirely, citing “previous work” without verification (Jiashuo2024ToG; chen2024pog).
@@
We observe significant misalignments between metric descriptions and their codebase implementations. For instance, although linhao2024rog, mavromatis2024gnn, and he2024g describe their primary metric as the ranking-based “Hits@1”, their official implementations entirely bypass ranking mechanisms. Instead, linhao2024rog and mavromatis2024gnn adopt a permissive criterion where any occurrence of the gold answer within the generated text is counted as correct. Similarly, the definition of “accuracy” fluctuates between strict Exact Match and loose substring matching across different studies (Jiashuo2024ToG; chen2024pog; jiang2023structgpt). Because many works reuse reported numbers under the assumption of comparability, these inconsistencies propagate, rendering cross-paper conclusions unreliable.
To address these issues, we adhere to the fully standardized and formally defined protocol below.
@@ 4.2 Evaluation Setup
Since KG-RAG models typically produce raw text sequences as output, we extract the final prediction set $\mathcal{P}_{q}$ by applying string partitioning and normalizing, following linhao2024rog. Details of this postprocessing step are provided in Appendix E. Without specific justification, all entities are represented by randomly assigned indices without textual labels (e.g., “Barack Obama” becomes 39) to ensure that models rely solely on knowledge from the KG rather than memorized surface forms.
@@ 5.1 Overall Performance
We evaluate six representative KG-RAG methods—RoG (linhao2024rog), G-Retriever (he2024g), GNN-RAG (mavromatis2024gnn), PoG (chen2024pog), StructGPT (jiang2023structgpt), and ToG (sun2023think)—on the benchmark datasets introduced in Section 3.
@@ 5.4 Influence of Entity Labeling
Figure 4: F1-Score comparison of models using Private ID, Entity ID, and Text Label representations on FB15k-237 (a-b) and Wikidata5m (c-d) benchmark under incomplete and complete KG settings.
@@ 5.5 Case Study
To better understand the limitations of current KG-RAG methods, we analyze representative failure cases from our benchmark. Table 4 shows two illustrative examples: (1) failure to retrieve relevant reasoning paths and (2) incorrect answer generation despite correct retrieval. A quantitative analysis is provided in Appendix J.
@@ 7 Limitation
This focus on Horn rules also offers a practical advantage: such rules are interpretable and easy to verify for plausibility, ensuring the reliability and consistency of the constructed questions. In contrast, more expressive rule types are challenging to mine reliably, and even SOTA rule mining methods often produce noisy or implausible rules. To maintain benchmark quality, we therefore focus on Horn rules in this paper. Extending to richer rule types is a promising direction for future work, specifically, incorporating advanced rule-learning methods like DRUM (sadeghian2019drum) and AnyBURL (meilicke2024anytime), or aggregating high-confidence rules from multiple algorithms, offers a principled way to expand coverage beyond the capabilities of any single system.

### v4

# What Breaks Knowledge Graph based RAG? Benchmarking and Empirical Insights into Reasoning under Incomplete Knowledge

@@ Abstract
Knowledge Graph-based Retrieval-Augmented Generation (KG-RAG) is an increasingly explored approach for combining the reasoning capabilities of large language models with the structured evidence of knowledge graphs. However, current evaluation practices fall short: existing benchmarks often include questions that can be directly answered using existing triples in KG, making it unclear whether models perform reasoning or simply retrieve answers directly. Moreover, inconsistent evaluation metrics and lenient answer matching criteria further obscure meaningful comparisons. In this work, we introduce a general method for constructing benchmarks and present BRINK (Benchmark for Reasoning under Incomplete Knowledge) to systematically assess KG-RAG methods under knowledge incompleteness. Our empirical results show that current KG-RAG methods have limited reasoning ability under missing knowledge, often rely on internal memorization, and exhibit varying degrees of generalization depending on their design.
What Breaks Knowledge Graph based RAG? Benchmarking and Empirical Insights into Reasoning under Incomplete Knowledge
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has become a widely adopted framework for enhancing large language models (LLMs) by incorporating external knowledge through a retrieve-then-generate paradigm (Khandelwal et al., 2020; Izacard and Grave, 2021; Borgeaud et al., 2022; Ram et al., 2023; Zhu et al., 2025b). By conditioning the generation on the retrieved documents, RAG enables LLMs to answer questions or perform tasks using more comprehensive and up-to-date knowledge than what is stored in their parameters. To improve retrieval accuracy, enable structured reasoning and support explanation, recent research has pivoted toward RAG methods based on Knowledge Graph (KG-RAG) (Han et al., 2024; Peng et al., 2024). Most of them directly use existing knowledge graphs (KGs) (Luo et al., 2024; Sun et al., 2024; Zhou et al., 2025b), while some of them construct and extend structured knowledge from unstructured documents (Fang et al., 2024). Such KG-RAG systems are expected to be well-suited for structured reasoning, where answers require synthesizing information from multiple connected facts.
Despite growing interest, current evaluation practices for KG-RAG fall short in two key ways. First, most existing benchmarks  (Yih et al., 2016; Talmor and Berant, 2018b) are constructed on top of complete KGs, where direct evidence supporting the answer is readily available. For example, given the question “Who is the brother of Justin Bieber?”, the KG contains the triple hasBrother(JustinBieber, JaxonBieber), allowing the system to answer the question without reasoning. However, real-world KGs are often incomplete, and answering such questions in practice may require reasoning over alternative paths, e.g., combining hasParent(JustinBieber, JeremyBieber) and hasChild(JeremyBieber, JaxonBieber) to infer the sibling relationship. As a result, current benchmarks do not assess whether KG-RAG methods can reason over missing knowledge or simply retrieve answers directly from explicit evidence.
@@
In this work, we introduce BRINK (Benchmark for Reasoning under Incomplete Knowledge). Constructed using a novel generalizable method, BRINK includes an evaluation protocol designed to systematically assess KG-RAG methods under knowledge incompleteness. Each question in our benchmark is constructed such that it cannot be answered using a single explicit triple. Instead, the answer must be inferred by reasoning over alternative paths in the KG. To construct this benchmark, we follow a two-step process: (1) we mine high-confidence logical rules from the KG using a rule mining algorithm, and (2) we generate natural language questions based on rule groundings, ensuring that the directly supporting triple is removed while the remaining KG still contains sufficient evidence to infer the answer.
Our empirical study on BRINK reveals several key limitations of current KG-RAG systems. First, most models struggle to recover answers when direct supporting facts are removed, highlighting their limited reasoning capacity. Second, training-based methods (e.g., RoG, GNN-RAG) show stronger robustness under KG incompleteness compared to non-trained systems. Third, we find that textual entity labels substantially boost performance, suggesting that LLMs rely heavily on internal memorization.
@@ 2.1 Knowledge Graph Question Answering (KGQA)
Approaches to KGQA can be broadly categorized into the following two types: (1) Semantic parsing-based methods Yih et al. (2016); Gu et al. (2021); Ye et al. (2021) translate questions into formal executable queries (e.g., SPARQL or logical forms), which are then executed over the KG to retrieve the answer entities. These methods offer high precision and interpretability, as the reasoning process is explicitly encoded. However, they face several challenges, including difficulty in understanding semantically and syntactically complex questions, handling diverse and compositional logical forms, and managing the large search space involved in parsing multi-relation queriesLan et al. (2021). (2) Embedding-based methods Yao et al. (2019); Baek et al. (2023), by contrast, encode questions and entities into a shared vector space and rank answer candidates based on embedding similarity. While these methods are end-to-end trainable and do not require annotated logical forms, they often struggle with multi-hop reasoning Qiu et al. (2020), lack interpretability Biswas et al. (2023), and exhibit high uncertainty in their predictions Zhu et al. (2024, 2025c, 2025a, 2025d).
@@ 2.2 KGQA with RAG
Recently, a new line of work has emerged that integrates large language models more tightly into the KG reasoning process. These KG-RAG methods go beyond previous KGQA approaches by coupling structured retrieval with generative reasoning capabilities of LLMs. RoG Luo et al. (2024) adopts a planning-retrieval-reasoning pipeline, where an LLM generates relation paths as plans, retrieves corresponding paths from the KG, and reasons over them for answer generation. G-Retriever He et al. (2024) retrieves subgraphs via Prize-Collecting Steiner Tree optimization and provides them to LLMs as soft prompts or text prompts for question answering. GNN-RAG Mavromatis and Karypis (2024) uses a GNN to select candidate answers and retrieves shortest paths to them, which are verbalized for LLM-based answer generation. PoG Chen et al. (2024) instead decomposes questions into subgoals and iteratively explores and self-corrects reasoning paths using a guidance-memory-reflection mechanism. ToG Sun et al. (2024) treats the LLM as an agent that interactively explores and aggregates evidence on KGs through iterative beam search, StructGPT Jiang et al. (2023) introduces an interface-based framework that enables LLMs to iteratively extract relevant evidence from structured data such as KGs and reason over the retrieved information to answer complex questions.
@@ 2.3 KGQA Benchmarks
Existing benchmarks such as WebQSP Yih et al. (2016), CWQ Talmor and Berant (2018a), and GrailQA Gu et al. (2021) are widely used to evaluate KGQA. However, during dataset construction, only question-answer pairs that yield a gold SPARQL answer on the reference KG are retained, while any unanswerable questions are discarded Yih et al. (2016); Talmor and Berant (2018a); Gu et al. (2021). This implicitly assumes that each question can be answered directly using an existing fact in the KG, overlooking the reality that KGs are often incomplete. To address this, recent works Xu et al. (2024); Zhou et al. (2025a) simulate KG incompleteness by either randomly deleting triples from the whole KG or removing those along the shortest path(s) between the question and answer entities. However, these approaches have a key limitation: they cannot ensure that sufficient knowledge remains in the KG to support answering each question. This can lead to misleading evaluations, as performance drops may stem from unanswerable questions rather than limitations in the model’s reasoning ability.
@@ 2.4 LLM Reasoning Evaluation
Evaluation of LLM reasoning typically covers both logical question answering and logical consistency, with benchmarks spanning deduction, multiple-choice, proof generation, and consistency constraints such as entailment, negation, and transitivity. Datasets like LogicBench Parmar et al. (2024), ProofWriter Tafjord et al. (2020), and LogicNLI Tian et al. (2021) are widely used for these tasks, but results indicate that LLMs still struggle with robust logical reasoning Cheng et al. (2025); He et al. (2025).
@@ 3.1 Rule Mining
To ensure that questions in our benchmark require reasoning rather than direct lookup, we first identify triples that are logically inferable from other facts. We achieve this by mining high-confidence Horn rules from the original KG using the AMIE3 algorithm (Lajus et al., 2020).
AMIE3 is a widely used rule mining system designed to operate efficiently over large-scale KGs. A logical rule discovered by AMIE3 has the following form (Horn rules Horn (1951)):
@@
We retain only rules with high confidence and sufficient support, filtering out noisy or spurious patterns. Specifically, we run AMIE3 with a confidence threshold of 0.3, a head coverage threshold of 0.1, and a maximum rule length of 4. AMIE3 incrementally generates candidate rules via breadth-first refinement Lajus et al. (2020) and evaluates them using confidence and head coverage; only those meeting the specified thresholds are retained. Additional details on the rule generation and filtering process are provided in Appendix A.
@@ 3.2 Dataset Generation
KGs typically exhibit a "long-tail" distribution, where a small number of entities participate in a disproportionately large number of triples, while the majority appear only infrequently  (Mohamed et al., 2020; Chen et al., 2023). This imbalance can cause many generated questions to share the same answer entity, leading to biased evaluation.
@@ 3.3 BRINK Overview
To support a systematic evaluation of reasoning under knowledge incompleteness, we construct BRINK based on three well established KGs: Family Sadeghian et al. (2019), FB15k-237 Toutanova and Chen (2015) and Wikidata5m Wang et al. (2021). These datasets differ in size, structure, and domain coverage, enabling evaluation across both synthetic and real-world settings.
@@
  • FB15k-237 is a widely used benchmark derived from Freebase Dong et al. (2014). The graph spans 14,541 entities and 237 predicates, covering real-world domains such as people, locations, and organizations.
  • Wikidata5m is a large-scale real-world KG constructed from Wikidata (Vrandečić and Krötzsch, 2014). It contains about 4.6 million entities, 822 relations, and over 20 million triples spanning diverse domains such as people, places, organizations, and events Wang et al. (2021).
@@
Table 2: Examples from BRINK datasets. Each instance includes a natural-language question, a topic entity (provided to the KG-RAG model), and the full set of correct answers. The red-highlighted answer denotes the hard answer—i.e., the one whose supporting triple has been removed in the incomplete KG setting. We also show the corresponding direct evidence (the deleted triple) and an alternative path derived from a mined rule that enables inference of the answer.
Datasets in BRINK.
@@
Table 3: Statistics of BRINK.
@@ 4.1 Standardization Gaps in Prior Work
Many studies rely on vague textual descriptions such as “whether the top-1 predicted answer is correct” for Hits@1 without specifying critical details (Luo et al., 2024; Mavromatis and Karypis, 2024; He et al., 2024). It is often unclear how LLM outputs are parsed, how “top-1” is defined in the absence of explicit ranking, and what constitutes correctness (e.g., exact match vs. substring inclusion). Some works omit definitions entirely, citing “previous work” without verification (Sun et al., 2024; Chen et al., 2024).
@@
We observe significant misalignments between metric descriptions and their codebase implementations. For instance, although Luo et al. (2024), Mavromatis and Karypis (2024), and He et al. (2024) describe their primary metric as the ranking-based “Hits@1”, their official implementations entirely bypass ranking mechanisms. Instead, Luo et al. (2024) and Mavromatis and Karypis (2024) adopt a permissive criterion where any occurrence of the gold answer within the generated text is counted as correct. Similarly, the definition of “accuracy” fluctuates between strict Exact Match and loose substring matching across different studies (Sun et al., 2024; Chen et al., 2024; Jiang et al., 2023). Because many works reuse reported numbers under the assumption of comparability, these inconsistencies propagate, rendering cross-paper conclusions unreliable.
To address these issues, we adhere to the fully standardized and formally defined protocol in BRINK.
@@ 4.2 Evaluation Setup
Since KG-RAG models typically produce raw text sequences as output, we extract the final prediction set $\mathcal{P}_{q}$ by applying string partitioning and normalizing, following Luo et al. (2024). Details of this postprocessing step are provided in Appendix E. Without specific justification, all entities are represented by randomly assigned indices without textual labels (e.g., “Barack Obama” becomes 39) to ensure that models rely solely on knowledge from the KG rather than memorized surface forms.
@@ 5.1 Overall Performance
We evaluate six representative KG-RAG methods—RoG (Luo et al., 2024), G-Retriever (He et al., 2024), GNN-RAG (Mavromatis and Karypis, 2024), PoG (Chen et al., 2024), StructGPT (Jiang et al., 2023), and ToG (Sun et al., 2023)—on the BRINK datasets introduced in Section 3.
@@ 5.4 Influence of Entity Labeling
Figure 4: F1-Score comparison of models using Private ID, Entity ID, and Text Label representations on FB15k-237 (a-b) and Wikidata5m (c-d) datasets under incomplete and complete KG settings.
@@ 5.5 Case Study
To better understand the limitations of current KG-RAG methods, we analyze representative failure cases from BRINK. Table 4 shows two illustrative examples: (1) failure to retrieve relevant reasoning paths and (2) incorrect answer generation despite correct retrieval. A quantitative analysis is provided in Appendix J.
@@ 7 Limitation
This focus on Horn rules also offers a practical advantage: such rules are interpretable and easy to verify for plausibility, ensuring the reliability and consistency of the constructed questions. In contrast, more expressive rule types are challenging to mine reliably, and even SOTA rule mining methods often produce noisy or implausible rules. To maintain benchmark quality, we therefore focus on Horn rules in this paper. Extending to richer rule types is a promising direction for future work, specifically, incorporating advanced rule-learning methods like DRUM (Sadeghian et al., 2019) and AnyBURL (meilicke2024anytime), or aggregating high-confidence rules from multiple algorithms, offers a principled way to expand coverage beyond the capabilities of any single system.